If the perimeter is 210 yards and the area is 2250 square yards, what are the dimensions?
Answer by checkley77(12569)

sORRY i DIDN'T NOTICE THE RECTANGLE IN THE 'Move to a better place box'.
XY=2250
X=2250/Y
2(2250/Y)+2Y=210
4500/Y+2Y=210
(4500+2Y*Y)/Y=210
4500+2Y^2=210Y
2Y^2-210Y+4500=0
2(Y^2-105+2250)=0
2(Y-75)(Y-30)=0
Y-75=0
Y=75
Y-30=0
Y=30
Y=75 THEN X=30
Y=30 THEN X=75
PROOF:
2*75+2*30=210
150+60=210
75*30=2250
2250=2250